OpenStudy (anonymous):

Find the product. (fg)^9 is it fg^9??..

mathslover (mathslover):

nope it will be : \[\large{f^9g^9}\]

OpenStudy (anonymous):

NO..

mathslover (mathslover):

Always remember that \[\large{(ab)^c=a^cb^c}\]

OpenStudy (anonymous):

Give power to both f and g..

mathslover (mathslover):

Got it @jonathanbae ?

OpenStudy (anonymous):

yep got it thanks

mathslover (mathslover):

*Nice to hear*

OpenStudy (usukidoll):

extremely simple (fg)^9 means writing f 9 times and writing g 9 times. distribute the 9 like this (f)^9(g)^9 =f^9g^9

mathslover (mathslover):

@UsukiDoll A request to you .. Please dont classify the question whether it is extremely simple or normal or extremely hard ..This may low down the user's confidence http://openstudy.com/code-of-conduct thanks!!!

OpenStudy (usukidoll):

ouch. sorry I'm new here. my bad

mathslover (mathslover):

No problem that is nice to see that you accepted it thanks and welcome to this site

OpenStudy (usukidoll):

:). thanks. I may need this site in the future. I'm going into Calculus III